# DaBiT: Depth and Blur informed Transformer for Joint Refocusing and Super-Resolution

### v1

@@ Abstract
In many real-world scenarios, recorded videos suffer from accidental focus blur, and while video deblurring methods exist, most specifically target motion blur. This paper introduces a framework optimised for the joint task of focal deblurring (refocusing) and video super-resolution (VSR). The proposed method employs novel map guided transformers, in addition to image propagation, to effectively leverage the continuous spatial variance of focal blur and restore the footage. We also introduce a flow re-focusing module to efficiently align relevant features between the blurry and sharp domains. Additionally, we propose a novel technique for generating synthetic focal blur data, broadening the model’s learning capabilities to include a wider array of content. We have made a new benchmark dataset, DAVIS-Blur, available. This dataset, a modified extension of the popular DAVIS video segmentation set, provides realistic out-of-focus blur degradations as well as the corresponding blur maps. Comprehensive experiments on DAVIS-Blur demonstrate the superiority of our approach. We achieve state-of-the-art results with an average PSNR performance over 1.9dB greater than comparable existing video restoration methods. Our source code will be made available at https://github.com/crispianm/DaBiT
@@ 1 Introduction
Videos captured in challenging circumstances, such as low light, or by less experienced shooters may exhibit blurriness, including motion blur and out-of-focus blur. Consequently, a deblurring process is necessary to enhance visual quality. However, this remains a challenging task, even for modern deep learning models, because of the rapid loss of information as the blur intensifies.
@@
While traditional methods employ deconvolution-based algorithms in order to sharpen input images, advances in deep learning techniques have enabled direct learning of a mapping from blurry to sharp. Recent contributions in this field, particularly those using convolutional neural networks (CNNs), have significantly improved upon the deconvolution-based algorithms, as demonstrated by image-based contributions such as DeBlurGAN [kupyn2019deblurgan], NAFNet [chen2022simple], and BIPNet [dudhane2021burst]. However, a significant challenge still exists in the form of video content; while these methods may produce visually pleasing results for individual frames, they often introduce temporal inconsistencies when applied to a video sequence.
@@
One explanation for this is that CNNs can struggle to capture long-range dependencies within video sequences; these are highly beneficial for handling complex image degradations like blur while retaining high-frequency details. The transformer architecture [dosovitskiy2020image] has emerged as a promising alternative for various computer vision tasks, due largely to its self-attention mechanism. The transformer’s ability to learn these global dependencies and accurately model long-range interactions makes it particularly well-suited for video deblurring, due to the inherent temporal dependencies of blur, as demonstrated by [liang2022rvrt, liang2024vrt]. While these methods are well-suited to removing camera shake and motion blur, scenarios exist where motion blur is desirable, and where errors in autofocus or camera operation also cause segments of the video to be out-of-focus unintentionally. In such cases, focal points will vary dynamically with the content, obscuring a large area of frames in the sequence. The spatio-temporal variance inherent in this effect makes the out-of-focus blur problem significantly more challenging to resolve than ’simple’ motion blur. Motivated by this, we introduce a Depth and Blur informed Transformer (DaBiT), which aims to learn the mapping from blurry to focused, employing pseudo ground truth blur and depth maps for inference.
The main contributions of this work are as follows:
  • The introduction of DaBiT, a novel transformer-based architecture integrating depth and blur maps, addressing the challenge of video defocus deblurring.
  • The adoption of a novel video inpainting approach to deblurring, focusing the model on out-of-focus sections.
  • The development of a novel synthetic blur generation process for realistic training data.
  • The integration of a video super-resolution component to handle high resolution content efficiently, enhancing the model’s performance.
@@ 2 Related Work
Video Deblurring. Deblurring aims to remove blur caused by incorrect focus, object motion, or camera shake from input videos. It can be accomplished either as individual images [zamir2021multi, abuolaim2020defocus, kong2023efficient, tsai2022stripformer, zamir2022restormer] or as sequences of images [hyun2015generalized, jin2018learning, zhong2023blur, kim2017dynamic, liang2024vrt, liang2022rvrt]. Importantly, only a handful of the individual image works aim to remove blur caused by incorrect focus [abuolaim2020defocus, son2021single, ruan2022learning, zamir2022restormer, zhao2022united, cui2023image, chen2023better].
@@
When methods trained to deblur individual images are applied to videos, temporal inconsistencies are prone to appear, due to small errors and inconsistencies in each image being restored differently. This necessitates the creation of specific network architectures for video deblurring. Recently, Zhong et allet@token. proposed a Blur Interpolation Transformer [zhong2023blur] based on Swin transformer blocks, that achieved competitive performance for motion deblurring. Shang et allet@token. attempted the joint task of frame interpolation and motion deblurring with VIDUE [shang2023joint], improving temporal consistency significantly for the challenging 16$\times$ interpolation task. To our knowledge, no work exists aiming to remove focal blur from entire video sequences.
@@
Video Super-Resolution. Video Super-Resolution (VSR) is a well studied task, with many significant recent research contributions. VSR is however challenging when compared to image super-resolution: not only by virtue of higher computational requirements, but also due to the challenge of maintaining temporal consistency across frames. In order to remedy this, two primary approaches have been proposed: sliding window-based and recurrent-based methods. While sliding window-based VSR methods [jo2018deep, li2020mucan, tian2020tdan, li2023simple] estimate the upscaled frames using adjacent frames in the pixel space, recurrent-based methods [huang2015bidirectional, liu2022learning, chan2021basicvsr, chan2022basicvsr++] propagate latent features sequentially to synthesise the HR target frame.
@@
No specific model architecture has yet dominated, and a variety of different model backbones have been utilised for this task. In 2023, Zhou et allet@token. proposed Upscale-A-Video [zhou2023upscale] introducing the concept of using text-guided diffusion for video upscaling, achieving a high perceptual quality. Most recently, Xu et allet@token. proposed VideoGigaGAN [xu2024videogigagan], balancing the inherent conflict of preserving high-frequency details whilst maintaining temporal consistency using their high-frequency shuffle component.
3 Synthetic Blurred Video Generation
Deep learning necessitates a substantial volume of data to train the model effectively. While unsupervised learning techniques exist, supervised learning-based methods still outperform them in terms of both accuracy and speed. However, obtaining paired datasets of blurred and clean sequences is time-consuming and content diversity can be limited. Hence, we introduce a new approach to generate synthetic datasets, and use it to create DAVIS-Blur, a modified version of the DAVIS [perazzi2016benchmark] video segmentation test set.
3.1 Focal Blur Simulation
In order for our model to learn the nature of focal blur, an accurate blur model is required.
We begin with an input video $I_{in}=\{I_{t}\in\mathbb{R}^{3\times H\times W}\}_{t=1}^{T}$ with $T$ frames. We then compute the depth maps $D=\{D_{t}\in\mathbb{R}^{1\times H\times W}\}_{t=1}^{T}$ for the input using DepthAnything [depthanything], due to its ability to predict temporally consistent depths at a level surpassing other state-of-the-art models such as MiDaS [ranftl_towards_2020] or ZoeDepth [bhat2023zoedepth]. Using the depth maps, the spatial variance of focal blur is able to be modelled. We begin by setting the focal point at a depth $f$, and defining a focal range $f_{r}$ centred on the focal point. We then apply gaussian blur kernels $G(x,y)=\frac{1}{2\pi\sigma^{2}}e^{-\frac{x^{2}+y^{2}}{2\sigma^{2}}}$, with $\sigma=5$, to the image in both directions along its depth axis, scaling symmetrically from zero inside the focal range to a maximum kernel size of $n_{max}\times n_{max}$. In this way, we blur the furthest depths from the focal point the most, creating a realistic focal blur effect.
During training, a random frame sequence from $I_{in}$ of length $l,l<T$ is taken, along with $r$ random global reference frames from elsewhere in the video. A random focal point ($f$) and focal range ($f_{r}$) are selected, along with a random focus rate $\frac{df}{dt}$, dictating the velocity at which the focal point will change in the sequence. A random maximum blur kernel size $n_{max}\in\{x\in\mathbb{Z}:3\leq x\leq 11,x\text{ is odd}\}$ is chosen, and the sequence and its global reference frames are blurred in a temporally variant manner, denoted as $\widetilde{I_{in}}$, simulating a focus shift. The spatial blur intensity information is recorded and used to create a ground truth blur map $B=\{B_{t}\in\mathbb{R}^{1\times H\times W}\}_{t=1}^{T}$, assisting our model during training. We also introduce data augmentation by means of a random horizontal flip and random reversal of sequence during the training. Finally, to enable super-resolution, we downsample the blurred sequence, along with its blur and depth maps, denoted as $\widetilde{I_{lr}}$, $B_{lr}$, and $D_{lr}$, respectively. Further details of this approach are provided in the supplementary material.
The variation in blur intensity and focus speed simulates a broad range of focus shifts. Some are representative of sequences which do not contain any additional global information (e.glet@token. $\frac{df}{dt}=0$, so $l$ is out-of-focus in the same spatio-temporal regions), and some which contain enough focused footage to be able to nearly reconstruct the whole sequence using only image propagation (e.glet@token. $f$ is varied completely across $\text{min}(D)$ and $\text{max}(D)$). This allows a given model to learn to deblur fully out-of-focus frames while not discarding the information contained in focused frames.
For creating our test set, we set $l=T$, $n_{max}=7$, $f_{r}=100$, and fix the initial focal point $f=0$ for all 50 sequences in the DAVIS [perazzi2016benchmark] test set. The focus rate is then set at $\text{max}(D)/l$, according to each sequence’s length. This simulates a complete focus shift from near to far, across the full length of the sequence. The large range of sequence lengths in the DAVIS set allow a wide range of focus shifts to be synthesised, thus posing a challenging task. The resulting new dataset, DAVIS-Blur, is released as a part of this work.
@@ 4 Methodology
Figure 1: The architecture of the DaBiT model.
@@
The proposed DaBiT framework is shown in Figure 1. The frames and pre-computed inputs to both the model and the flow restoration module are downsampled to allow footage of high resolution to be processed faster, also adding flexibility in other use cases, such as video editing.
Given a low-resolution, blurry input video $I_{lr}$ with $T$ frames and corresponding blur and depth maps $B_{lr}$ and $D_{lr}$, our aim is to remove the focal blur in a spatio-temporally consistent manner and jointly upscale the video to high resolution. During inference, the blur maps can be estimated using a combination of the depth maps and a wavelet transform. By taking high-frequency regions of the frame found with the transform as regions of focus, we take the sum of the wavelet components for each depth, and use the inverse of this as a map of the blurred regions. These maps are needed to employ our inpainting-based approach to this task, under the assumption that there are other spatio-temporally aligned regions in the video which exhibit less or no blur. This takes the form of integrating blur masks, allowing our model to identify the regions of focus in a given input, and it reduces the difficulty of the task greatly.
@@
To generate the blur mask sequence $M=\{M_{t}\in\mathbb{R}^{1\times H\times W}\}_{t=1}^{T}$, we begin with the blur map sequence $B_{lr}$, where the intensity of the grayscale image corresponds to the amount of blur present in the spatial region. The maps are binarised by setting all but the lowest value equal to one, trivially distinguishing regions of focus from those containing blur, and delivering the resulting masks, $M_{lr}$. Separately, we also employ RAFT [teed2020raft] to estimate the blurry input frames’ flows, $\widetilde{F}=\{\widetilde{F_{t}}\in\mathbb{R}^{2\times H\times W}\}_{t=1}^{T-1}$. The flows are passed to the refocusing module (subsection 4.1) to obtain in-focus flows $F_{lr}$. After refocusing, $\widetilde{I_{lr}}$ and $M_{lr}$ are passed to a non-learnable image propagation module [zhou2023propainter], which warps flows $F$ to recover details present at different times in the sequence, using forward and backward consistency error [gao2020flow, xu2019deep]. It returns updated input frames $\widetilde{I_{lr}}^{\prime}$ as well as the corresponding updated masks $M_{lr}^{\prime}$, dilating the spatio-temporal focused area. The blur maps are left unaltered in the event that propagation delivers blurred or temporally inaccurate content.
@@
$I_{lr}^{\prime},M_{lr}^{\prime},D_{lr}$, and $B_{lr}$ are then passed to an encoder, which compresses the inputs into a latent representation, enabling a significant reduction in compute for the map guided transformers. Next, the latent features are passed into the map-guided sparse transformer blocks, processed, and decoded into frames. Finally, the deblurred frames are passed to an upscaler, which achieves 2$\times$ super-resolution, returning $I_{out}$. While the potential is not limited, upscaling was only tested to 2$\times$ due to memory constraints.
The specifics of each component are detailed below.
@@ 4.1 Flow Refocusing Module
In order for the image propagation to be most effective, optical flows are required to be as close to ground truth as possible. Similarly to ProPainter [zhou2023propainter], which employs a pre-trained recurrent flow completion network for inpainting objects from flows, we design a pre-trained network for aiding in the refocusing of flows. During inference, the optical flows estimated by RAFT [teed2020raft] are between blurry, low-resolution frames. Errors in the flow have the potential to propagate into the reconstructed footage as artefacts and temporal inconsistencies. To reduce the likelihood of this, we train a separate network specifically to refocus the degraded flow, due to the distinct difference in task between refocusing flows and refocusing video sequences. We follow previous works [chan2021basicvsr, chan2022basicvsr++, zhou2023propainter] in choosing a recurrent network architecture, due to its efficiency and performance relative to a sliding window approach.
Figure 2: The architecture of the Flow Refocusing module.
@@
The architecture of this module can be seen in Figure 2. During training, we create samples in the same manner detailed in subsection 3.1, and use binarised blur maps as masks for the flows. We downsample the inputs by a factor of eight for increased speed, and pass the masked flows, binary masks, and blurry frames to an encoder. Following [chan2022basicvsr++], we perform deformable alignment based on deformable convolution [dai2017deformable, zhu2019deformable], propagating in-focus regions from other flows and lessening the burden of refocusing of the module. We then decode the features into refocused flows, which are used in the image propagation module to reduce the burden on the transformer.
4.2 DaBiT

### v3

# DaBiT: Depth and Blur informed Transformer for Video Focal Deblurring

@@ Abstract
In many real-world scenarios, recorded videos suffer from accidental focus blur, and while video deblurring methods exist, most specifically target motion blur or spatial-invariant blur. This paper introduces a framework optimized for the as yet unattempted task of video focal deblurring (refocusing). The proposed method employs novel map-guided transformers, in addition to image propagation, to effectively leverage the continuous spatial variance of focal blur and restore the footage. We also introduce a flow re-focusing module designed to efficiently align relevant features between blurry and sharp domains. Additionally, we propose a novel technique for generating synthetic focal blur data, broadening the model’s learning capabilities and robustness to include a wider array of content. We have made a new benchmark dataset, DAVIS-Blur, available. This dataset, a modified extension of the popular DAVIS video segmentation set, provides realistic focal blur degradations as well as the corresponding blur maps. Comprehensive experiments demonstrate the superiority of our approach. We achieve state-of-the-art results with an average PSNR performance over 1.9dB greater than comparable existing video restoration methods. Our source code and the developed databases will be made available at https://github.com/crispianm/DaBiT
@@ 1 Introduction
Videos captured in challenging circumstances, such as low light, or by less experienced shooters may exhibit blurriness, including motion blur and out-of-focus blur. Specifically, in scenarios with extremely high noise, it is difficult for a camera operator or autofocus algorithm to determine if a subject is in focus. This often results in out-of-focus footage after a denoising post-process is applied. Consequently, a focal deblurring process is necessary to recover high-frequency details, perform focus correction, and improve overall visual quality. However, this remains a challenging task, even with modern deep learning models, because of the loss of information as the blur intensifies.
@@
While traditional methods employ deconvolution-based algorithms in order to sharpen input images [2], advances in deep learning techniques have enabled direct learning of a mapping from blurry to sharp. Recent contributions in this field, particularly those using convolutional neural networks (CNNs), have significantly improved upon the deconvolution-based algorithms, as demonstrated by image-based contributions such as DeBlurGAN [24], NAFNet [6], and BIPNet [14]. However, a significant challenge still exists in the form of video content; while these methods may produce visually pleasing results for individual frames, they often introduce noticeable temporal inconsistencies when applied to a video sequence.
@@
One explanation for this is that CNNs may struggle to capture long-range dependencies within video sequences; these are highly beneficial for handling complex image degradations like blurs while retaining high-frequency details. The transformer architecture [13] has emerged as a promising alternative for various computer vision tasks, largely due to its self-attention mechanism. The transformer’s ability to learn these global dependencies and accurately model long-range interactions makes it particularly well suited for video focal deblurring, due to the inherent temporal dependencies of blur, as demonstrated by [29, 28]. While these methods have been shown to be suitable for removing camera shake and motion blur, scenarios exist where motion blur is desirable, and some undesirable focal blur is present. In such cases, focal points will vary dynamically with the content, obscuring a large amount of frames in the sequence. The spatio-temporal variance inherent in this effect makes the video focal blur problem significantly more challenging to resolve than ‘simple’ motion blur.
Figure 1: The architecture of the DaBiT model. Green highlights represent the temporally variant regions of the flows and input frames which are out of focus.
In this context, this paper proposes a Depth and Blur informed Transformer (DaBiT), which aims to learn the mapping from blurry to focused. As a novel transformer-based architecture, DaBiT (illustrated in Figure 1) employs depth and blur maps for video refocusing and deblurring. It also utilizes a novel spatial selection approach which reduces the computational burden, enabling recovery of information from out-of-focus areas. The DaBiT framework further incorporates a video super-resolution module, enabling flexible input resolutions. To optimize the proposed model, a large synthetic blurred video dataset is procedurally generated through randomized realistic focal blur simulation. The same generation method is also employed to create a test set, DAVIS-Blur, based on the video segmentation dataset, DAVIS [36]. The main contributions of this work are as follows:
  • 1)
    The proposed DaBiT model is the first to attempt the task of focus-related video deblurring.
  • 2)
    We are the first to employ a spatial selection method to recover the areas which are out of focus, using depth and blur maps for added guidance.
  • 3)
    We propose a novel focal blur simulator and generate both training data and test dataset ‘DAVIS-Blur’ for optimizing and validating focus-related video deblurring methods. Both will be released publicly along with the DaBiT model.
The proposed DaBiT model was evaluated on our DAVIS-Blur dataset and the low-light video BVI-RLV dataset [30], and compared against six existing video completion, super-resolution, and refocusing models. The results demonstrated the superior performance of DaBiT, with an average PSNR gain of more than 1.9 dB.
@@ 2.1 Video Deblurring
Deblurring aims to remove blur caused by incorrect focus, object motion, or camera shake from input videos. It can be accomplished either as individual images [51, 1, 23, 43, 50] or as sequences of images [18, 19, 55, 21, 28, 29]. Importantly, only a handful of the single image based models aim to remove blur caused by incorrect focus [1, 40, 38, 50, 54, 10, 8].
@@
When methods trained to deblur individual images are applied to videos, temporal inconsistencies are prone to appear, due to small errors and inconsistencies in each image being restored differently. This necessitates the creation of specific network architectures for video deblurring. Recently, Zhong et al. proposed a Blur Interpolation Transformer [55] based on Swin transformer blocks, that achieved competitive performance for motion deblurring. Shang et al. attempted the joint task of frame interpolation and motion deblurring with VIDUE [39], improving temporal consistency significantly for the challenging 16$\times$ interpolation task. To our knowledge, no work exists aiming to remove focal blur from entire video sequences.
2.2 Video Super-Resolution
@@
Another ill-posed problem closely related to deblurring is super-resolution. Video Super-Resolution (VSR) is a well studied task, with many significant recent research contributions. It is however considered simpler than refocusing because the degradation in VSR is usually uniform across the frame. VSR is challenging when compared to image super-resolution; not only by virtue of higher computational requirements, but also due to the challenge of maintaining temporal consistency across frames. In order to remedy this, two primary approaches have been proposed: sliding window-based and recurrent-based methods. While sliding window-based VSR methods [20, 26, 42, 25] estimate the upscaled frames using adjacent frames in the pixel space, recurrent-based methods [17, 31, 4, 5] propagate latent features sequentially to synthesize the HR target frame. BasicVSR++ has emerged as a prominent tool for this task, due in large part to the effectiveness of its proposed second-order grid propagation and flow-guided deformable alignment [5].
@@
No specific model architecture has yet dominated, and a variety of different model backbones have been utilized for this task. In 2023, Zhou et al. proposed Upscale-A-Video [57] introducing the concept of using text-guided diffusion for video upscaling, achieving a high perceptual quality. Most recently, Xu et al. proposed VideoGigaGAN [47], balancing the inherent conflict of preserving high-frequency details whilst maintaining temporal coherence through a high-frequency shuffling mechanism.
Figure 2: The architecture of the Flow Refocusing module. Green highlights represent regions of the flows which are out of focus.
@@ 3 Methodology
3.1 Overview of DaBiT
@@
The proposed DaBiT framework is shown in Figure 1. The frames and pre-computed inputs to both the model and the flow restoration module are downsampled to reduce the complexity and memory required, enabling footage of high resolution to be processed faster, also adding flexibility in other use cases, such as video editing.
Given a low-resolution, blurry input video $\widetilde{I_{in}}$ with $T$ frames and corresponding blur and depth maps $B_{in}$ and $D_{in}$, our aim is to remove the focal blur in a spatio-temporally consistent manner and jointly upscale the video to high resolution.
@@
The process starts with generating $B_{in}$ and the corresponding binary mask $M_{in}$ (subsection 3.2), and employs RAFT [41] to estimate the blurry input frames’ flows, $\widetilde{F}=\{\widetilde{F_{t}}\in\mathbb{R}^{2\times H\times W}\}_{t=1}^{T-1}$. The flows are passed to the refocusing module (subsection 3.3) to obtain in-focus flows $F_{in}$. After refocusing, $\widetilde{I}_{in}$ and $M_{in}$ are passed to a non-learnable image propagation module [56], which warps flows $F_{in}$ to recover details present at different times in the sequence, using forward and backward consistency error [15, 46]. It returns updated input frames $\widetilde{I}_{in}^{\prime}$ as well as the corresponding updated masks $M_{in}^{\prime}$, dilating the spatio-temporal focused area. The blur maps are left unaltered in the event that propagation delivers blurred or temporally inaccurate content.
@@
$I_{in}^{\prime},M_{in}^{\prime},D_{in}$, and $B_{in}$ are then passed to the DaBiT network (subsection 3.4), where an encoder compresses the inputs into a latent representation, enabling a significant reduction in compute for the map guided transformers. Next, the latent features are passed into the map-guided sparse transformer blocks, processed, and decoded into frames. Finally, the deblurred frames are passed to an upscaler, which achieves 2$\times$ super-resolution, returning $I_{out}$. The addition of super-resolution at this stage gives the model the ability to add extra details lost due to blur. To minimize excess parameters, we limit super-resolution to 2$\times$, as it is experimentally found to offer an excellent trade-off between performance and complexity when compared to 4$\times$ SR.
By integrating blur and depth maps into a super-resolution transformer framework, DaBiT not only corrects out-of-focus regions but also enhances video quality, making it particularly effective for high-resolution video restoration. The specifics of each component are detailed below.
3.2 Blur map $B_{in}$ and binary mask $M_{in}$ generation
During inference, the blur maps can be estimated using a combination of the depth maps and a wavelet transform. By taking high-frequency regions of the frame found with the transform as regions of focus, we take the sum of the wavelet components for each depth, and use the inverse of this as a map of the blurred regions. These maps are needed to employ our spatial selection approach to this task, under the assumption that there are other spatio-temporally aligned regions in the video which exhibit less or no blur. This takes the form of integrating blur masks, allowing our model to identify the regions of focus in a given input, and it reduces the difficulty of the task greatly.
To generate the binary mask sequence $M=\{M_{t}\in\mathbb{R}^{1\times H\times W}\}_{t=1}^{T}$, we begin with the blur map sequence $B_{in}$, where the intensity of the grayscale image corresponds to the amount of blur present in the spatial region. The maps are binarized by setting all but the lowest value equal to one, trivially distinguishing regions of focus from those containing blur, and delivering the resulting masks, $M_{in}$.
@@ 3.3 Flow Refocusing Module
In order for the image propagation to be most effective, optical flows are required to be as close to ground truth as possible. Similarly to other models which employ pre-trained flow completion networks for inpainting objects from flows, we design a recurrent pre-trained network for aiding in the refocusing of flows. During inference, the optical flows estimated by RAFT [41] are between blurry, low-resolution frames. Errors in the flow have the potential to propagate into the reconstructed footage as artifacts and temporal inconsistencies. To reduce the likelihood of this, we train a separate network specifically to refocus the degraded flow, due to the distinct difference in task between refocusing flows and refocusing video sequences. We follow previous works [4, 5, 56] in choosing a recurrent network architecture, due to its efficiency and performance relative to a sliding window approach.
@@
The architecture of this module can be seen in Figure 2. During training, we create samples in the same manner detailed in subsection 4.1, and use binarized blur maps as masks for the flows. We downsample the inputs by a factor of eight for increased speed, and pass the masked flows, binary masks, and blurry frames to an encoder. Following [5], we perform deformable alignment based on deformable convolution [11, 58], propagating in-focus regions from other flows and lessening the burden of refocusing of the module. We then decode the features into refocused flows, which are used in the image propagation module to reduce the burden on the transformer.
3.4 Map Guided Transformer